# Supervision, Writing – review and editing

### v1

@@ Abstract
The growing availability of longitudinal Magnetic Resonance Imaging (MRI) datasets has facilitated Artificial Intelligence (AI)-driven modeling of disease progression, making it possible to predict future medical scans for individual patients. However, despite significant advancements in AI, current methods continue to face challenges including achieving patient-specific individualization, ensuring spatiotemporal consistency, efficiently utilizing longitudinal data, and managing the substantial memory demands of 3D scans. To address these challenges, we propose Brain Latent Progression (BrLP), a novel spatiotemporal model designed to predict individual-level disease progression in 3D brain MRIs. The key contributions in BrLP are fourfold: (i) it operates in a small latent space, mitigating the computational challenges posed by high-dimensional imaging data; (ii) it explicitly integrates subject metadata to enhance the individualization of predictions; (iii) it incorporates prior knowledge of disease dynamics through an auxiliary model, facilitating the integration of longitudinal data; and (iv) it introduces the Latent Average Stabilization (LAS) algorithm, which (a) enforces spatiotemporal consistency in the predicted progression at inference time and (b) allows us to derive a measure of the uncertainty for the prediction. We train and evaluate BrLP on 11,730 T1-weighted (T1w) brain MRIs from 2,805 subjects and validate its generalizability on an external test set comprising 2,257 MRIs from 962 subjects. Our experiments compare BrLP-generated MRI scans with real follow-up MRIs, demonstrating state-of-the-art accuracy compared to existing methods. The code is publicly available at: https://github.com/LemuelPuglisi/BrLP.
@@ 1 Introduction
    Individualization. Disease progression is influenced by various individual factors, including demographic and clinical data. To improve prediction accuracy, models must incorporate and leverage subject-specific metadata.
@@
To solve these challenges, we introduce Brain Latent Progression (BrLP), a novel individual-based spatiotemporal model capable of predicting disease progression on 3D brain MRIs at the individual level. BrLP offers several key contributions to address the outlined challenges. First, we propose combining a Latent Diffusion Model (LDM) (Rombach et al., 2022) with a ControlNet (Zhang et al., 2023) to generate individualized brain MRIs conditioned on available subject data, addressing challenge 1. Second, we integrate prior knowledge of disease progression by employing an auxiliary model that infers volumetric changes in different brain regions, enabling the use of longitudinal data when available and addressing challenge 2. Third, we introduce Latent Average Stabilization (LAS), a technique to improve spatiotemporal consistency in the predicted progression, addressing challenge 3. Fourth, we utilize latent representations of brain MRIs to reduce memory demands for processing 3D scans, addressing challenge 4. Finally, we demonstrate how LAS can be used to derive a measure of uncertainty in the predictions, which could serve as a reliability metric in clinical applications.
@@
This work extends our MICCAI 2024 conference article (Puglisi et al., 2024) in several ways: (1) we enrich the ablation study by analyzing the hyperparameter for the LAS algorithm; (2) we test BrLP on an external dataset to evaluate its generalization capabilities; (3) we evaluate the impact of the cognitive status as a conditioning variable; (4) we introduce a mechanism to quantify the uncertainty of predictions within the BrLP framework and provide statistical analysis to support our findings; and (5) we showcase an example of a potential clinical application of BrLP for patient selection in clinical trials.
@@ 2 Related work
In this section, we describe prior work on spatiotemporal modeling, which can be broadly categorized into population-based and individual-based models. We then focus on individual-based models and review how previous studies have leveraged the latest advances in generative AI to produce synthetic scans for this task.
2.1 Population-based models
@@
Population-based models estimate an average trajectory of disease progression in a high-dimensional space using data from a population of affected subjects. The resulting average trajectory provides an interpretable insight into the disease dynamics. A key challenge in these models is mapping individual subjects onto a common disease timeline, which cannot correspond directly to chronological age due to variations in age at disease onset and progression rate. For example, in (Schiratti et al., 2015), the authors define a general spatiotemporal model using Riemannian geometry. This method estimates the average disease trajectory as a geodesic on a Riemannian manifold and considers individual trajectories as curves parallel to the average geodesic. The method uses time reparameterization to map individuals to the shared disease timeline. The Riemannian formulation allows the application of the method to high-dimensional data, such as full brain MRIs or 3D shapes, but requires the definition of a suitable Riemannian metric. In a related study (Sauty and Durrleman, 2022b) the Riemannian metric is learned from the data, reducing the inductive bias of the model at the cost of an extra computational burden. A recent work (Sauty and Durrleman, 2022a) bridges the gap between traditional population-based models and novel deep learning techniques, using a Variational Autoencoder (VAE) to encode brain MRIs into a latent space, in which it defines a linear mixed effect model to learn the average disease trajectory from the population. Similarly, in (Chadebec et al., 2022), the authors use a VAE to project images into a latent space and fit a generative model for disease progression with a fully variational approach. These approaches can predict individual disease progression by adjusting the parameters of an average disease trajectory to align with a subject’s available longitudinal data. Although this solution is effective for scalar biomarkers, it becomes overly restrictive when applied to high-dimensional data, where progression patterns can display significant variability.
2.2 Individual-based models
@@
Individual-based models, also known as simulators (Ravi et al., 2022), aim to predict changes in high-dimensional data (e.g., a full MRI scan or 3D shape) for a given subject over a specified period and under specific conditions (e.g., in the presence of neurodegenerative disease). By operating at the individual level, these models can use each patient’s age as a temporal axis to track disease progression. Generally, individual-based models offer greater flexibility compared to population-based approaches, though this flexibility often comes at the cost of reduced interpretability of the underlying disease dynamics. Most of the existing individual-based techniques leverage advances in deep generative methods to estimate the conditional distribution of future medical images given the input data. The most popular models used for this task include Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), VAEs (Kingma, 2013), Normalizing Flows (NFs) (Papamakarios et al., 2021), and most recently, diffusion models (Ho et al., 2020). The next sections will describe these methods in greater detail.
@@ 2.2.1 Generative adversarial networks
GANs employ two competing neural networks — a generator that produces synthetic data and a discriminator that evaluates authenticity — to create increasingly realistic artificial outputs through iterative adversarial training. An example of this approach is proposed in (Xia et al., 2021). Here, the authors used a GAN-based method to simulate subject-specific brain aging trajectories, conditioned on the presence of AD, by minimizing adversarial and reconstruction losses without using longitudinal data. However, this method has only been applied to 2D slices, lacking a full 3D implementation. Among the first methods to model disease progression in 3D brain MRIs, the seminal work (Ravi et al., 2022) introduces 4D-DaniNet, a generative model that exploits adversarial learning to provide individualized predictions of brain MRIs. 4D-DaniNet incorporates prior knowledge of disease progression by embedding biological constraints into the loss function during training. To address memory limitations, 4D-DaniNet synthesizes 2D slices, which are subsequently reassembled into a 3D volume using a super-resolution module. However, this approach fails to capture inter-slice dependencies, which are critical for modeling inherently three-dimensional phenomena, highlighting the need for methodologies that operate directly on 3D images. In (Jung et al., 2021), the authors present ADESyn, a conditional GAN for synthesizing 3D brain MRIs across different stages of AD. The model generates 2D slices using an attention-based generator conditioned on disease progression and ensures 3D spatial consistency through integrated 2D and 3D discriminators. Instead of using the subject age as the temporal axis, ADESyn employs a disease condition score ranging from 0 (healthy) to 1 (AD), limiting its ability to distinguish changes in the early and late stages of AD. Lastly, CounterSynth (Pombo et al., 2023) is a GAN-based counterfactual synthesis method that can simulate various conditions within a brain MRI, including aging and disease progression. Rather than modeling the entire 3D brain MRI, CounterSynth generates a diffeomorphic transformation that warps the input image to reflect specified covariates (e.g., changes at a target age under a given condition). This technique preserves anatomical accuracy and minimizes artificial artifacts in the generated images. However, it is limited to modeling only structural changes.
@@ 2.2.2 Variational autoencoders
VAEs are deep generative models that learn to map data to and from a structured probabilistic latent space, enabling both compression and generation. In (He et al., 2024), the authors propose a double-encoder conditional VAE for predicting future MRI scans based on baseline MRI and subject metadata. The model also facilitates disease classification by estimating the posterior distribution when both baseline and follow-up MRIs are provided. Although demonstrated using 3D brain MRIs, the experiments are performed at low resolution (2 $2\text{ mm}^{3}$) to mitigate computational costs. Furthermore, the model does not address other challenges, such as the utilization of longitudinal data and enforcing the spatiotemporal consistency of the predicted progression.
@@ 2.2.3 Normalizing flows
NFs use sequences of invertible transformations to convert simple distributions into complex ones while maintaining exact likelihood evaluation. Based on this, (Wilms et al., 2022) introduces a bidirectional NF model that links brain morphology to age, potentially incorporating additional variables, such as disease diagnosis. This approach enables the generation of follow-up images at a target age and the estimation of brain age from a given image. Their model leverages NFs and, similar to CounterSynth (Pombo et al., 2023), utilizes diffeomorphic deformations to represent structural changes. While this bidirectional framework holds promise, it cannot utilize longitudinal data if available, and it lacks mechanisms to enforce spatiotemporal consistency.
@@ 2.2.4 Diffusion models
In recent years, diffusion models have emerged as a major breakthrough in generative AI, gaining significant attention for their state-of-the-art image synthesis capabilities and stable training processes, which avoid the challenges associated with adversarial approaches. A Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is a deep generative model with two Markovian processes: forward diffusion and reverse diffusion. In the forward process, Gaussian noise is incrementally added to the original image $x_{0}$ over $T$ steps. At each step $t$, noise is introduced to the current image $x_{t-1}$ by sampling from a Gaussian transition probability defined as $q(x_{t}\mid x_{t-1})\coloneqq\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ follows a variance schedule. If $T$ is sufficiently large, $x_{T}$ will converge to pure Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$. The reverse diffusion process aims to revert each diffusion step, allowing the generation of an image from the target distribution starting from pure noise $x_{T}$. The reverse transition probability has a Gaussian closed form, $q(x_{t-1}\mid x_{t},x_{0})=\mathcal{N}(x_{t-1}\mid\tilde{\mu}(x_{0},x_{t}),\tilde{\beta}_{t})$, conditioned on the real image $x_{0}$. As $x_{0}$ is not available during generation, a neural network is trained to approximate $\mu_{\theta}(x_{t},t)\approx\tilde{\mu}(x_{0},x_{t})$. Following the work proposed in (Ho et al., 2020), it is possible to reparameterize the mean in terms of $x_{t}$ and a noise term $\epsilon$, and then use a neural network to predict the noise $\epsilon_{\theta}(x_{t},t)\approx\epsilon$, optimized with the following objective:
@@
Diffusion models have been successfully applied to 3D medical image synthesis, showing promising results (Pinaya et al., 2022). In the context of spatiotemporal modeling, the Sequence-Aware Diffusion Model (SADM) (Yoon et al., 2023) marks a significant advancement by allowing for the generation of longitudinal brain scans through autoregressive sampling informed by sequential MRI data. While SADM improves upon previous spatiotemporal methods by leveraging longitudinal data, it has several limitations, including poor temporal control, high memory requirements, inability to incorporate subject metadata, and a lack of guaranteed spatiotemporal consistency in its predictions. Other works have applied diffusion-based approaches to spatiotemporal modeling, but they are either limited to 2D slices of the brain (Litrico et al., 2024) or focused on different diseased organs (Kim and Ye, 2022). Our proposed method, BrLP, addresses the limitations of existing approaches by harnessing the latest advances in diffusion modeling to deliver an individual-based framework that integrates subject-specific metadata, effectively utilizes available longitudinal data, enforces spatiotemporal consistency, and handles high-dimensional imaging data without sacrificing modeling capacity. A key innovation of our approach is its treatment of prior knowledge: unlike 4D-DaniNet, which incorporates prior knowledge of disease progression by embedding biological constraints into the loss function during training, BrLP takes a more direct approach by integrating this prior knowledge directly into the image generation process itself.
Figure 1: The overview of BrLP training and inference process. The training process outputs an autoencoder (A) that maps 3D brain MRIs into small latent representations; (B) an LDM able to generate latent representations according to subject-specific and progression-related covariates; (C) a ControlNet, able to constrain the LDM’s generation process to a subject’s brain. During inference (E), progression-related variables at the target age are first predicted by an auxiliary model (D). These predictions, combined with subject-specific variables and the baseline MRI, condition the generation of the latent representations corresponding to the predicted brain at the target age. Finally, the LAS algorithm (F) repeats this process $m$ times and averages the obtained latent representations before decoding the result into the 3D MRI space.

### v2

@@ Abstract
The growing availability of longitudinal Magnetic Resonance Imaging (MRI) datasets has facilitated Artificial Intelligence (AI)-driven modeling of disease progression, making it possible to predict future medical scans for individual patients. However, despite significant advancements in AI, current methods continue to face challenges including achieving patient-specific individualization, ensuring spatiotemporal consistency, efficiently utilizing longitudinal data, and managing the substantial memory demands of 3D scans. To address these challenges, we propose Brain Latent Progression (BrLP), a novel spatiotemporal model designed to predict individual-level disease progression in 3D brain MRIs. The key contributions in BrLP are fourfold: (i) it operates in a small latent space, mitigating the computational challenges posed by high-dimensional imaging data; (ii) it explicitly integrates subject metadata to enhance the individualization of predictions; (iii) it incorporates prior knowledge of disease dynamics through an auxiliary model, facilitating the integration of longitudinal data; and (iv) it introduces the Latent Average Stabilization (LAS) algorithm, which (a) enforces spatiotemporal consistency in the predicted progression at inference time and (b) allows us to derive a measure of the uncertainty for the prediction at the global and voxel level. We train and evaluate BrLP on 11,730 T1-weighted (T1w) brain MRIs from 2,805 subjects and validate its generalizability on an external test set comprising 2,257 MRIs from 962 subjects. Our experiments compare BrLP-generated MRI scans with real follow-up MRIs, demonstrating state-of-the-art accuracy compared to existing methods. The code is publicly available at: https://github.com/LemuelPuglisi/BrLP.
@@ 1 Introduction
    Individualization. Disease progression is influenced by various individual factors, including demographic and clinical variables. To improve prediction accuracy, models must incorporate and leverage subject-specific metadata.
@@
To solve these challenges, we introduce Brain Latent Progression (BrLP), a novel individual-based spatiotemporal model capable of predicting disease progression on 3D brain MRIs at the individual level. BrLP offers several key contributions to address the outlined challenges. First, we propose combining a Latent Diffusion Model (LDM) (Rombach et al., 2022) with a ControlNet (Zhang et al., 2023) to generate individualized brain MRIs conditioned on available subject data, addressing challenge 1. Second, we integrate prior knowledge of disease progression by employing an auxiliary model that infers volumetric changes in different brain regions, enabling the use of longitudinal data when available and addressing challenge 2. Third, we introduce Latent Average Stabilization (LAS), a technique to improve spatiotemporal consistency in the predicted progression, addressing challenge 3. Fourth, we utilize latent representations of brain MRIs to reduce memory demands for processing 3D scans, addressing challenge 4. Finally, we demonstrate how LAS can be used to derive a measure of uncertainty in the predictions, both at the global and voxel level, which could serve as a reliability metric in clinical applications.
@@
This work extends our MICCAI 2024 conference article (Puglisi et al., 2024) in several ways: (1) we enrich the ablation study by analyzing the hyperparameter for the LAS algorithm; (2) we test BrLP on an external dataset to evaluate its generalization capabilities; (3) we evaluate the impact of the cognitive status as a conditioning variable; (4) we introduce a mechanism to quantify the uncertainty of predictions, both at the global and voxel level, within the BrLP framework and provide statistical analysis to support our findings; and (5) we showcase an example of a potential clinical application of BrLP for patient selection in clinical trials.
@@ 2 Related work
In this section, we review prior work on spatiotemporal disease progression modeling, which can be broadly categorized into population-based and individual-based approaches. We then focus on the latter, examining how recent studies have leveraged advances in generative AI to produce synthetic scans for individual-level disease progression. Finally, we discuss the limitations of existing approaches and how our proposed method addresses these challenges.
2.1 Population-based progression modeling
@@
Population-based progression modeling involves learning an average disease trajectory from longitudinal patient data, often using methods such as mixed-effects models (Schiratti et al., 2017). Within this framework, individual predictions are obtained by modeling deviations from the average trajectory, allowing for a degree of personalization. A central challenge in such models is mapping individual subjects along a unified disease timeline, as chronological age fails to account for critical inter-patient variations in both age at disease onset and rate of progression. For example, in (Schiratti et al., 2015), the authors define a general spatiotemporal model using Riemannian geometry. This method estimates the average disease trajectory as a geodesic on a Riemannian manifold and considers individual trajectories as curves parallel to the average geodesic. The method uses time reparameterization to map individuals to the shared disease timeline. While the Riemannian formulation is highly generalizable, it requires the definition of a suitable Riemannian metric. In a related study (Sauty and Durrleman, 2022b), the Riemannian metric is learned from the data, reducing the inductive bias of the model at the cost of an extra computational burden. A recent work (Sauty and Durrleman, 2022a) bridges the gap between the previous methods and novel deep learning techniques, using a Variational Autoencoder (VAE) to encode brain MRIs into a latent space, in which it defines a linear mixed effect model to learn the average disease trajectory from the population. Similarly, in (Chadebec et al., 2022), the authors use a VAE to project images into a latent space and fit a generative model for disease progression with a fully variational approach. While population-based models provide advantages—such as an interpretable representation of average disease dynamics—they can become overly constrained when applied to high-dimensional data, where individual progression patterns vary significantly.
2.2 Individual-based progression modeling
@@
Individual-based progression modeling is a more flexible framework in which models are directly trained to map current individual observations to their corresponding states at future timepoints. By operating at the individual-level, these models can use the subject’s chronological age as the temporal axis, allowing for a better personalized modeling of disease trajectories. While individual-based models offer greater flexibility than population-based approaches, this often comes at the expense of interpretability regarding the underlying disease mechanisms. Many of these models leverage recent advances in deep generative techniques, making them particularly well-suited for handling high-dimensional data such as full MRI scans or 3D anatomical shapes. The most popular frameworks used for this task include Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), VAEs (Kingma, 2013), Normalizing Flows (NFs) (Papamakarios et al., 2021), and most recently, diffusion models (Ho et al., 2020). The next sections will describe these methods in greater detail.
@@ 2.2.1 Generative adversarial networks
GANs employ two competing neural networks — a generator that produces synthetic data and a discriminator that evaluates authenticity — to create increasingly realistic artificial outputs through iterative adversarial training. An example of this approach is proposed in (Xia et al., 2021). Here, the authors used a GAN-based method to simulate subject-specific brain aging trajectories, conditioned on the presence of AD, by minimizing adversarial and reconstruction losses without using longitudinal data. Among the first methods to model disease progression in 3D brain MRIs, the seminal work (Ravi et al., 2022) introduces 4D-DaniNet, a generative model that exploits adversarial learning to provide individualized predictions of brain MRIs. To address memory limitations, 4D-DaniNet synthesizes 2D slices, which are subsequently reassembled into a 3D volume using a super-resolution module. 4D-DaniNet incorporates prior knowledge of disease progression by embedding biological constraints into the training loss function. In contrast, our proposed BrLP model introduces a key innovation by integrating prior knowledge directly into the image generation process, guiding synthesis at inference time. In (Jung et al., 2021), the authors present ADESyn, a conditional GAN for synthesizing 3D brain MRIs across different stages of AD. The model generates 2D slices using an attention-based generator conditioned on disease progression and ensures 3D spatial consistency through integrated 2D and 3D discriminators. Instead of using the subject age as the temporal axis, ADESyn employs a disease condition score ranging from 0 (healthy) to 1 (AD), limiting its ability to distinguish changes in the early and late stages of AD. Lastly, CounterSynth (Pombo et al., 2023) is a GAN-based counterfactual synthesis method that can simulate various conditions within a brain MRI, including aging and disease progression. Rather than modeling the entire 3D brain MRI, CounterSynth generates a diffeomorphic transformation that warps the input image to reflect specified covariates (e.g., changes at a target age under a given condition). This technique preserves anatomical accuracy and minimizes artificial artifacts in the generated images.
@@ 2.2.2 Variational autoencoders
VAEs are deep generative models that learn to map data to and from a structured probabilistic latent space, enabling both compression and generation. In (He et al., 2024), the authors propose a double-encoder conditional VAE for predicting future MRI scans based on baseline MRI and subject metadata. The model also facilitates disease classification by estimating the posterior distribution when both baseline and follow-up MRIs are provided.
@@ 2.2.3 Normalizing flows
NFs use sequences of invertible transformations to convert simple distributions into complex ones while maintaining exact likelihood evaluation. Based on this, (Wilms et al., 2022) introduces a bidirectional NF model that links brain morphology to age, potentially incorporating additional variables, such as disease diagnosis. This approach enables the generation of follow-up images at a target age and the estimation of brain age from a given image. Their model leverages NFs and, similar to CounterSynth (Pombo et al., 2023), utilizes diffeomorphic deformations to represent structural changes.
@@ 2.2.4 Diffusion models
In recent years, diffusion models have emerged as a major breakthrough in generative AI, gaining significant attention for their state-of-the-art image synthesis capabilities and stable training processes, which avoid the challenges associated with adversarial approaches. A detailed overview of the diffusion model framework is provided in Section 3. These models have been successfully applied to 3D medical image synthesis, showing promising results (Pinaya et al., 2022). In the context of spatiotemporal modeling, the Sequence-Aware Diffusion Model (SADM) (Yoon et al., 2023) marks a significant advancement by allowing for the generation of longitudinal brain scans through autoregressive sampling informed by sequential MRI data. Other works have applied diffusion-based approaches to spatiotemporal modeling, but they are either limited to 2D slices of the brain (Litrico et al., 2024) or focused on different diseased organs (Kim and Ye, 2022). Building on recent advancements, our proposed BrLP method also uses a latent diffusion framework extended with a ControlNet (Zhang et al., 2023) to more effectively model spatiotemporal dynamics, as detailed in the methodology section.
2.3 Limitations of the existing methods
In this section, we relate existing methodologies to the four key challenges outlined in Section 1. The first challenge pertains to the level of individualization a model can achieve. Predictions from population-based progression models are not truly individualized, as the underlying model remains grounded in population-level patterns rather than being fully adapted to each patient’s unique data. In contrast, individual-based models offer greater flexibility and personalization. Nevertheless, their potential is often underutilized, as many existing approaches (Jung et al., 2021; Wilms et al., 2022; Yoon et al., 2023) do not incorporate subject-specific metadata (e.g., demographic or diagnostic information) that could enhance individual-level predictions. The second challenge involves the effective exploitation of longitudinal data. Among the reviewed methods, only the work by (Yoon et al., 2023) explicitly leverages temporal sequences during inference, whereas the other approaches disregard this valuable information. The third challenge concerns spatiotemporal consistency. Some earlier GAN-based approaches address this by enforcing identity preservation (Xia et al., 2021; Jung et al., 2021) or incorporating biological constraints (Ravi et al., 2022). However, these solutions have seen diminishing attention, with more recent approaches (Pombo et al., 2023; He et al., 2024; Wilms et al., 2022; Yoon et al., 2023) omitting explicit mechanisms to ensure temporally coherent and biologically plausible progression. The fourth and final challenge is the memory demand associated with processing 3D medical images. Several methods address this by operating on 2D slices and subsequently stacking them to form a 3D volume (Xia et al., 2021; Ravi et al., 2022; Jung et al., 2021). While computationally efficient, this strategy fails to capture inter-slice dependencies, which are critical for modeling inherently three-dimensional phenomena, highlighting the need for methodologies that operate directly on 3D images. Other approaches rely on lightweight deformation fields to model temporal progression as spatial transformations of a baseline image (Wilms et al., 2022; Pombo et al., 2023). While deformation fields can preserve anatomical structures, they are inherently constrained to reshaping pre-existing anatomical content. As a result, they cannot generate new structures (e.g., white matter hypointensities in T1w images) not present in the initial scan. Another strategy is to rely on low-resolution volumes (He et al., 2024), which risks omitting clinically relevant anatomical details.
Our proposed method, BrLP, addresses the limitations of existing approaches by harnessing the latest advances in diffusion modeling to deliver an individual-based framework that integrates subject-specific metadata, effectively utilizes available longitudinal data, enforces spatiotemporal consistency, and handles high-dimensional imaging data without sacrificing modeling capacity.
3 Background - Diffusion models
@@
A Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is a deep generative model with two Markovian processes: forward diffusion and reverse diffusion. In the forward process, Gaussian noise is incrementally added to the original image $x_{0}$ over $T$ steps. At each step $t$, noise is introduced to the current image $x_{t-1}$ by sampling from a Gaussian transition probability defined as $q(x_{t}\mid x_{t-1})\coloneqq\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, where $\beta_{t}$ follows a variance schedule. If $T$ is sufficiently large, $x_{T}$ will converge to pure Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$. The reverse diffusion process aims to revert each diffusion step, allowing the generation of an image from the target distribution starting from pure noise $x_{T}$. The reverse transition probability has a Gaussian closed form, $q(x_{t-1}\mid x_{t},x_{0})=\mathcal{N}(x_{t-1}\mid\tilde{\mu}(x_{0},x_{t}),\tilde{\beta}_{t})$, conditioned on the real image $x_{0}$. As $x_{0}$ is not available during generation, a neural network is trained to approximate $\mu_{\theta}(x_{t},t)\approx\tilde{\mu}(x_{0},x_{t})$. Following the work proposed in (Ho et al., 2020), it is possible to reparameterize the mean in terms of $x_{t}$ and a noise term $\epsilon$, and then use a neural network to predict the noise $\epsilon_{\theta}(x_{t},t)\approx\epsilon$, optimized with the following objective:
@@
Figure 1: The overview of BrLP training and inference process. The training process outputs an autoencoder (A) that maps 3D brain MRIs into small latent representations; an LDM (B) able to generate latent representations according to subject-specific and progression-related covariates; a ControlNet (C), able to constrain the LDM’s generation process to a subject’s brain. During inference (E), progression-related variables at the target age are first predicted by an auxiliary model (D). These predictions, combined with subject-specific variables and the baseline MRI, condition the generation of the latent representations corresponding to the predicted brain at the target age. Finally, the LAS algorithm (F) repeats this process $m$ times and averages the obtained latent representations before decoding the result into the 3D MRI space.